2019 economy.
Consumption expenditure (C) 9 000, Investment (I) 6 500, Government spending (G) 7 000, Exports (X) 1 800, Imports (Z) 2 400, Depreciation 700, Foreign payment to the rest of the world 300, Foreign payment from the rest of the world 250

1. Calculate the value of the country’s GDE (Gross Domestic Expenditure)
2. Compute the value for the country’s GDP (Gross Domestic Product) at market price
3. Determine the value of the country’s NNI (Net National Income) at market price
4. If it is predicted that the GDP will increase to 22 000 in 2020, calculate the growth rate between 2019 and 2020
1
Expert's answer
2020-12-03T10:47:39-0500

y=C+I+G(X-M)

Y=9000+6500+7000+(1800-2400)

Y=22500-600

Y=21900

Gross domestic expenditure is 21900

b,

GDP at market price=final consumption+G+Net domestic capital formation

9000+6500+7000-600

21900

c,

NNI at market price

GDP at market price -depreciation+net factor income from abroad

21900-700+250=21450

d growth rate

22000-21900=100

100/22000*100

0.5%